# Seyed Ali Farokh

### v1

@@ Abstract
This paper presents our submissions to the Iranian division of the Text-dependent Speaker Verification Challenge (TdSV) 2024. TdSV aims to determine if a specific phrase was spoken by a target speaker. We developed two independent subsystems based on pre-trained models: For phrase verification, a phrase classifier rejected incorrect phrases, while for speaker verification, a pre-trained ResNet293 with domain adaptation extracted speaker embeddings for computing cosine similarity scores. In addition, we evaluated Whisper-PMFA, a pre-trained ASR model adapted for speaker verification, and found that, although it outperforms randomly initialized ResNets, it falls short of the performance of pre-trained ResNets, highlighting the importance of large-scale pre-training. The results also demonstrate that achieving competitive performance on TdSV without joint modeling of speaker and text is possible. Our best system achieved a MinDCF of 0.0358 on the evaluation subset and won the challenge.
Index Terms: Speaker Recognition, Transfer Learning
@@ 1 Introduction
Speaker verification (SV) is the process of confirming a person’s identity through their voice. This verification uses one or more enrollment utterances along with a test utterance and can be either text-dependent (TdSV) or text-independent (TiSV). In a TiSV system, the phonetic content of the utterance is unrestricted, allowing for any spoken content. In contrast, TdSV systems verify both the speaker’s identity and the specific phrase spoken. These systems can be either phrase-dependent, where a fixed set of phrases is pre-defined by the system, or phrase-independent, allowing users to customize their own phrases [1]. With the rising demand for voice-based authentication, TdSV has recently gained more attention, as the phonetic content of utterances can be used as passphrases [2], adding an extra layer of security to voice-based access control systems.
Recent advancements in TiSV have largely focused on improving cascaded systems, which typically consist of a front-end for extracting fixed-size speaker-specific features and a backend for computing similarity scores between speaker embeddings [2]. For the backend, cosine similarity and probabilistic linear discriminant analysis (PLDA) [3] are widely used. The front-end is usually a deep speaker classification network that comprises a frame-level subnetwork for initial feature extraction, a temporal pooling layer for computing fixed-size speaker embeddings from variable-length feature sequences, and a final subnetwork for classification. ResNets [4] and ECAPA-TDNNs [5] are among the most widely used architectures for front-ends, and the most common aggregation techniques include temporal statistics pooling [6] and variants of attentive pooling [7, 8]. Also, advanced softmax variants such as additive margin softmax (AM-Softmax) [9] loss and additive angular margin softmax (AAM-Softmax) [10] loss have been proposed to achieve more discriminative speaker embedding spaces. In addition, pre-trained automatic speech recognition (ASR) models have recently been adapted for speaker verification, showing promising results [11, 12, 13].
TiSV architectures can be effectively adapted for TdSV applications by joint modeling of linguistic and speaker information within a single network. One approach is to use multi-task learning to disentangle speaker and text information, often using two distinct subnetworks for speaker classification and phoneme (or phrase) classification [14, 15]. Further, Liu et al. proposed a phoneme-aware attentive pooling [16] method that incorporates frame-level phoneme posteriors into attentive pooling, enhancing the model’s ability to leverage phonetic information.
@@
This paper presents our system submitted to Task 1 of The Iranian division of The Text-dependent Speaker Verification (TdSV) Challenge 2024111Challenge website: https://tdsvc.github.io [17]. In this challenge, model enrollment is done using three enrollment utterances, and each trial consists of a test utterance and a model identifier. Trials fall into one of the following categories:
@@
  • Target Correct (TC): The test utterance belongs to the claimed model, and the speaker utters the correct phrase.
@@
  • Target Wrong (TW): The test utterance belongs to the claimed model, but the speaker utters an incorrect phrase.
@@
  • Imposter Correct (IC): The speaker’s identity does not match the claimed model, but the speaker utters the correct phrase.
  • Imposter Wrong (IW): The speaker’s identity does not match the claimed model, and the speaker utters an incorrect phrase.
Proposed systems should independently process each trial and generate a score that integrates both speaker and phrase verification scores. The generated scores should only accept TC trials. Also, note that IW trials were excluded from the challenge dataset because they can be easily rejected by a phrase verification system and do not provide valuable insights about a system’s overall performance.
@@
Task 1 of TdSV 2024 is phrase-dependent and uses a fixed set of ten phrases—five in Persian and five in English—for creating enrollment and test recordings. To increase the challenge, some test utterances in TW trials were drawn from free-text recordings.
@@
The primary evaluation metric adopted by TdSV 2024 is the normalized minimum Detection Cost Function (MinDCF), as defined in NIST SRE 2008 as a weighted sum of miss and false error probabilities, with $P_{target}=0.01$, $C_{FalseAlarm}=1$, and $C_{Miss}=10$. Additionally, the Equal Error Rate (EER) metric will be reported.
@@
Unlike most previous studies that adopted joint modeling of speaker and text, our proposed system comprises two independent subsystems for phrase and speaker verification, allowing for the use of various pre-trained models for each subsystem. For phrase verification, we fine-tuned a pre-trained cross-lingual speech representation model for bilingual speech recognition in Persian and English, followed by a further fine-tuning stage for phrase classification. This classifier was used to simply reject TW trials. Similarly, we developed several speaker embedding extractors based on pre-trained ResNets and Whisper [18] for our speaker verification system. After rejecting incorrect phrases using our phrase classifier, we obtained final verification scores by computing cosine similarity between test and enrollment embeddings. Moreover, we used score normalization to enhance the performance further. The experimental results indicated the effectiveness of the proposed method, and our top-performing model ranked first on the challenge leaderboard.
@@
2 Datasets
The DeepMine dataset [19, 20] is the primary source of the training and evaluation data for TdSV 2024. It was collected through crowd-sourcing, and while all participants were native Persian speakers, most contributed to the English portion of the dataset as well. The official TdSV 2024 data for Task 1 includes three subsets: training, development, and evaluation. The training subset consists of 183,431 utterances from 1,620 speakers. Among the utterances, 31,738 are free-text, while the rest were drawn from a fixed set of ten phrases comprising five Persian and five English phrases. The development and evaluation subsets are intended solely for system evaluation and contain 117,348 and 6,464,241 trials, respectively. During evaluation, model enrollment is conducted using three recordings of a specific phrase, and each trial includes a test utterance and a model identifier. The development set is provided to participants for evaluation and parameter tuning before submitting results to the official leaderboard, whereas the evaluation subset is used for the official evaluation of the challenge. In addition to the DeepMine dataset, participants are also allowed to use the following datasets:
  • VoxCeleb 1&2 [21, 22] are two large-scale datasets collected from YouTube videos. They contain over one million recordings from 7,205 celebrities. In this work, only VoxCeleb 1 was used, which includes over 100,000 utterances from 1,251 speakers.
  • LibriSpeech [23] is a standard ASR corpus in US English that comprises approximately 1,000 hours of speech from 2,338 speakers. We only used the train-clean-100 subset of this dataset to train our phrase verification system, which contains about 100 hours of speech.
  • Common Voice [24] is a multilingual speech dataset created from contributions by volunteers worldwide. For this challenge, teams are restricted to using the Persian (Farsi) subset, which contains approximately 363 hours of validated speech from 4,148 speakers [^2] Common Voice 18.0, released on 6/19/2024. To prepare this subset for training our speaker verification systems, we excluded speakers with fewer than 30 recordings. From the remaining speakers with more than 650 recordings, we randomly selected 650 utterances per speaker, resulting in a final dataset with 125,017 utterances from 813 speakers.
@@ 2.1 Data augmentation
We did not use any augmentation methods in our phrase verification system. However, following the previous successful studies on speaker verification [25], we adopted SoX-based  [^3] https://sourceforge.net/projects/sox speed perturbation by factors of 0.9 and 1.1 to triple the number of speakers during training, followed by an on-the-fly implementation of the following augmentations with a probability of 0.6:
  • Noise addition with MUSAN [26] dataset
  • RIR reverberation with RIRs [27] dataset
  • Gain augment
@@ 3 Phrase verification system
The phrase classifier is an 11-class model trained with standard softmax. The first ten classes correspond to the set of phrases in the challenge, and the final class represents free text (or ”none of the above”). This classifier is built on XLSR [28], a pre-trained cross-lingual speech representation model trained by solving a self-supervised contrastive task, proven to be effective in low-resource languages compared to traditional feature extraction methods. This model takes a raw waveform as input and produces a sequence of features.
Moreover, to improve the model’s ability to extract linguistic features from Persian and English inputs, we initially fine-tuned the XLSR for bilingual speech recognition in Persian and English. During this phase, 30% of the training subsets of Common Voice Farsi and LibriSpeech (train-clean-100) were used, and the model was trained using CTC loss [29] for 40 epochs, with an initial learning rate of 0.001 and an effective batch size of 32.
Finally, for training the classifier, an attention-based pooling layer was added to the fine-tuned XLSR to compute utterance-level feature vectors from frame-level representations $h_{t}$ ($t=1,...,T$):
@@
where $e_{t}$ and $\alpha_{t}$ are the attention score and weight, respectively. $\tilde{h}$ refers to the utterance-level feature vector, which is finally fed to a fully connected layer with ReLU activation, followed by a linear classifier. The network was trained using Cross-Entropy loss function for one epoch on the entire training samples of the challenge dataset, with a learning rate of 0.0005 and an effective batch size of 64.
Table 1: Hyper-parameters used across different training stages. BS and LR denote batch size and learning rate.
System Full Training Domain Adaptation
Epoch BS LR Epoch BS LR
S2 - - - 15 32 3e-4
S3 100 64 1e-3 15 32 3e-4
S4 15 64 1e-3 7 28 5e-5
S5 15 64 1e-3 7 28 5e-5
@@ 4 Speaker verification system
Our primary goal in Task 1 of TdSV 2024 was to leverage the power of pre-trained models to overcome the challenges of in-domain data scarcity and mitigate the computational cost of training randomly initialized models. We explored two directions for developing speaker verification models and compared their results. In the first approach, we trained several ResNet-based models, which have been widely used as a standard architecture in speaker verification. In the second approach, we studied the performance of pre-trained ASR models adapted for the speaker verification task. More specifically, we employed the Whisper-PMFA [30] method, which involves fine-tuning a pre-trained Whisper model for speaker recognition.
@@ 4.1 Training protocol
  • Full training (T1): In this stage, models were trained on a combination of out-of-domain data (Common Voice Farsi and VoxCeleb 1) and in-domain (DeepMine) data, totaling 3,684 speakers, to learn robust and generalizable speaker embeddings across different datasets. pre-trained ResNets did not undergo this stage. During this phase, 300 frames of each sample input were randomly selected in each epoch to prevent overfitting and accelerate the training. Moreover, all augmentation methods explained in Section 2.1 were applied. We employed the widely used AAM-Softmax [10] loss with the subcenter method and the Inter-TopK penalty [31] to train our models, with a constant margin and scale of 0.2 and 32, respectively.
  • Domain adaptation (T2): We fine-tuned our models with in-domain data after full training to bridge the domain gap and enhance the performance. During this stage, augmentation methods and the Inter-TopK penalty were removed to avoid domain mismatch, and the number of randomly selected frames was increased from 300 to 600. Also, fine-tuning was done with smaller learning rate values to maintain the generalization abilities of the models.
All models were optimized using SGD with a momentum of 0.9 and a weight decay of 1e-4. We also utilized an exponential decay scheduler with a minimum learning rate of 5e-5 for T1 and 1e-6 for T2. Other training hyper-parameters are listed in Table 1. Note that gradient accumulation was used to achieve the target effective batch size when GPU memory was limited. The dimensionality of speaker embeddings was set to 256 across all models. All experiments were conducted on a single Nvidia RTX 3090 GPU using the WeSpeaker toolkit [32, 33].
@@ 4.2 ResNet
ResNet [4] is a widely used architecture for speaker recognition that has performed excellently in previous speaker verification challenges [25]. Consequently, many open-source implementations and pre-trained models have been publicly released based on this architecture. Trained on large-scale datasets like VoxCeleb 1&2, these pre-trained models can provide a robust starting point for training speaker recognition models on other datasets by improving their generalization and speeding up the convergence.
During the challenge period, we submitted three systems based on bottleneck-block-based ResNet, all of which adopted the temporal statistics pooling [6] mechanism to aggregate variable-length sequence features to utterance-level embeddings. The first system (S1) was a pre-trained ResNet34 without domain adaptation, while the second one (S2) was a pre-trained ResNet293 that underwent domain adaptation. Finally, we applied both training stages to a randomly initialized ResNet152 to obtain our last ResNet-based system (S3).
Table 2: Phrase verification results on TC-vs-TW trials
Subset MinDCF0.01 EER(%)
development 0.0000 0.00
evaluation 0.0003 0.01
Table 3: Results on the development and evaluation sets
System Architecture Training Stages Development Evaluation
MinDCF0.01 EER(%) MinDCF0.01 EER(%)
S1 ResNet34 0.0614 1.3938 0.0784 1.7390
S2 ResNet293 T2 0.0225 0.8733 0.0376 1.1080
S3 ResNet152 T1 + T2 0.0191 0.6757 0.0764 2.3444
S4 Whisper-PMFA T1 + T2 0.0163 0.6121 0.0584 2.0410
S5 Whisper-PMFA T1 + T2 0.0161 0.6126 0.0583 2.0445
[…]

### v2

# Memory-Efficient Training for Text-Dependent SV with Independent Pre-trained Models

@@ Abstract
This paper presents our submission to the Iranian division of the Text-Dependent Speaker Verification Challenge (TdSV) 2024. Conventional TdSV approaches typically jointly model speaker and linguistic features, requiring unsegmented inputs during training and incurring high computational costs. Additionally, these methods often fine-tune large-scale pre-trained speaker embedding models on the target domain dataset, which may compromise the pre-trained models’ original ability to capture speaker-specific characteristics. To overcome these limitations, we employ a TdSV system that utilizes two pre-trained models independently and demonstrate that, by leveraging pre-trained models with targeted domain adaptation, competitive results can be achieved while avoiding the substantial computational costs associated with joint fine-tuning on unsegmented inputs in conventional approaches. Our best system reached a MinDCF of 0.0358 on the evaluation subset and secured first place in the challenge.
Keywords: Text-dependent Speaker Verification, Speaker Verification, Memory-efficient Training, Pre-trained Models, Transfer Learning
@@ 1 Introduction
Speaker verification (SV) is the task of confirming an individual’s identity based on their voice. It involves comparing one or more enrollment utterances with a test utterance and can be performed in either a text-independent (TiSV) or text-dependent (TdSV) setting. In TiSV, the phonetic content of the utterances is unrestricted, and only the speaker’s identity is verified, whereas in TdSV, the system verifies both the speaker’s identity and the specific phrase spoken. With the development of various neural network architectures (Xie et al., 2019; Desplanques et al., 2020; Zeinali et al., 2019b; Snyder et al., 2018), loss functions (Xiang et al., 2019; Zhang and Koishida, 2017; Wang et al., 2018; Deng et al., 2019), and pooling methods (Snyder et al., 2018; India et al., 2019; Zhu et al., 2018), TiSV has seen considerable improvement in recent years, whereas TdSV has remained relatively underexplored. TdSV systems can be either phrase-dependent (i.e., shared passphrases), where a fixed set of phrases is predefined by the system, or phrase-independent (i.e., user-defined passphrases), allowing users to customize their phrases (Hossein et al., 2024). With the growing demand for voice-based authentication, TdSV has gained increasing attention, as the phonetic content can be used as passphrases (Tu et al., 2022), adding an extra layer of security to voice-based access control systems.
@@
This paper presents our system submitted to Task 1 of the Text-dependent Speaker Verification Challenge 2024111Challenge website: https://tdsvc.github.io (Zeinali et al., 2025), which aimed to encourage participants to explore novel approaches for TdSV. The challenge was organized into two divisions: an international one, which included two subtasks focusing on shared and user-defined passphrases, and an Iranian division, which mirrored Task 1 of the worldwide challenge but specifically emphasized developing methods with limited GPU resources. In this challenge, model enrollment is done using three enrollment utterances, and each trial consists of a test utterance and a model identifier. Speaker verification trials fall into one of the following categories:
@@
  • Target Correct (TC): The speaker matches the claimed model and utters the correct phrase.
@@
  • Target Wrong (TW): The speaker matches the claimed model but utters an incorrect phrase.
@@
  • Impostor Correct (IC): The speaker does not match the claimed model but utters the correct phrase.
  • Impostor Wrong (IW): The speaker does not match the claimed model and utters an incorrect phrase. This category was excluded from the current year’s challenge, as it does not pose sufficient difficulty for contemporary models.
@@
In the context of TdSV, proposed systems are required to integrate both speaker and phrase verification scores and accept only TC trials [^2] For Text-independent Speaker Verification (TiSV), the task definition differs: both TC and TW trials are accepted.. Task 1 is phrase-dependent, employing a fixed set of ten phrases (five in Persian and five in English) for enrollment and testing. Additionally, to enhance the complexity of the challenge, some test utterances in TW trials were sourced from free-text recordings.
@@
The primary evaluation metric adopted by TdSV 2024 is the normalized minimum Detection Cost Function (MinDCF), as defined in NIST SRE 2008 as a weighted sum of miss and false error probabilities, with $P_{target}=0.01$, $C_{FalseAlarm}=1$, and $C_{Miss}=10$. The Equal Error Rate (EER) will also be reported as a secondary performance measure.
Previous successful approaches to TdSV typically jointly model speaker characteristics and the linguistic content of utterances. For instance, Liu et al. (2021) proposed a phoneme-aware attentive pooling method that incorporates frame-level phoneme posteriors into attentive pooling, improving the model’s ability to utilize phonetic information effectively. Also, some studies have employed supervised multi-task learning to jointly learn speaker and linguistic features for further improvement (Yang et al., 2020; Han et al., 2021).
However, joint speaker and phrase modeling has some drawbacks compared to independent modeling. First, model development becomes more complex than developing the system based on independent phrase and speaker embedding models. Additionally, since phrase modeling requires attending to an entire utterance, inputs cannot be chunked during training, requiring variable-length inputs to be zero-padded. This issue substantially increases GPU memory requirements, particularly for recent transformer-based models, due to their quadratic time and memory complexity (Vaswani et al., 2017).
Furthermore, as demonstrated in this work, pre-trained speaker embedding models are highly effective at extracting speaker-related features while disregarding other information in input utterances. However, when subjected to multi-task fine-tuning, these models are prone to lose their initial ability to extract speaker-related features, allocating capacity to learning linguistic content instead. This shift reduces their effectiveness, especially when in-domain data for multi-task fine-tuning is limited.
@@
Motivated by these challenges, we leverage the full capacity of pre-trained models and develop a TdSV system based on independent pre-trained models for phrase and speaker verification. For phrase verification, we fine-tune a pre-trained cross-lingual speech representation model for bilingual automatic speech recognition (ASR) in Persian and English, followed by a further fine-tuning stage for phrase classification. This classifier is used to reject incorrect phrases. Similarly, we develop several speaker embedding extractors based on pre-trained ResNets and Whisper (Radford et al., 2023) for our speaker verification system. After rejecting incorrect phrases using the phrase classifier, final verification scores are obtained by computing cosine similarity between test and enrollment embeddings.
Experimental results demonstrate that with well-designed fine-tuning stages, our TdSV system built on independently pre-trained models can achieve performance comparable to systems that jointly model speaker-related and linguistic information while using only a single Nvidia RTX 3090 GPU. This strategy substantially lowers GPU memory requirements and, consequently, reduces computational costs compared to the multi-GPU setups typically employed for training speaker recognition models (Zheng et al., 2023). Our best system secured first place in the Iranian division of the challenge and outperformed the third-place team in the international division (Zeinali et al., 2025).
@@
2 Challenge Datasets
The DeepMine dataset (Zeinali et al., 2018, 2019a) is the primary source of the training and evaluation data for TdSV 2024. It was collected through crowd-sourcing, and while all participants were native Persian speakers, most contributed to the English portion of the dataset as well. The official TdSV 2024 data for Task 1 includes three subsets: training, development, and evaluation. The training subset consists of 183,431 utterances from 1,620 speakers. Among the utterances, 31,738 are free-text, while the rest were drawn from a fixed set of ten phrases comprising five Persian and five English phrases. The development and evaluation subsets are intended solely for system evaluation and contain 117,348 and 6,464,241 trials, respectively. During evaluation, model enrollment is conducted using three recordings of a specific phrase, and each trial includes a test utterance and a model identifier. The development set is provided to participants for evaluation and parameter tuning before submitting results to the official leaderboard. The evaluation subset is used for the official evaluation of the challenge. In addition to the DeepMine dataset, participants are also allowed to use the following datasets:
  • VoxCeleb 1&2 (Nagrani et al., 2017; Chung et al., 2018) are two large-scale datasets collected from YouTube videos, which contain over one million recordings from 7,205 celebrities. In this work, due to resource constraints, only VoxCeleb 1 was used, which includes over 100,000 utterances from 1,251 speakers.
  • LibriSpeech (Panayotov et al., 2015) is a standard ASR corpus in US English that comprises approximately 1,000 hours of speech from 2,338 speakers. We only used the train-clean-100 subset of this dataset to train our phrase verification system, which contains about 100 hours of speech.
  • Common Voice (Ardila et al., 2020) is a multilingual speech dataset created from contributions of volunteers from worldwide. For this challenge, teams are restricted to using the Persian (Farsi) subset, which contains approximately 363 hours of validated speech from 4,148 speakers [^3] Common Voice 18.0, released on 6/19/2024. To prepare this subset for training our speaker verification systems, we excluded speakers with fewer than 30 recordings. From the remaining speakers with more than 650 recordings, we randomly selected 650 utterances per speaker, resulting in a final dataset with 125,017 utterances from 813 speakers.
@@ 2.1 Data Augmentation
We did not use any augmentation methods in our phrase verification system. However, following the previous successful studies on speaker verification (Chen et al., 2022; Zheng et al., 2023), we adopted SoX-based speed perturbation by factors of 0.9 and 1.1 to triple the number of speakers during training, followed by an on-the-fly implementation of the following augmentations, each applied with a probability of 0.6: noise addition using the MUSAN dataset (Snyder et al., 2015), reverberation using RIRs dataset (Ko et al., 2017), and gain augmentation.
@@ 3 Phrase Verification System
The phrase classifier is an 11-class model trained with standard softmax. The first ten classes correspond to the set of phrases in the challenge, and the final class represents free text (or “none of the above”). This classifier is built on XLSR [^4] Facebook/wav2vec2-xls-r-300m (Conneau et al., 2021), a pre-trained cross-lingual speech representation model trained by solving a self-supervised contrastive task, proven to be effective in low-resource languages compared to traditional feature extraction methods. This model takes a raw waveform as input and produces a sequence of features.
Moreover, to improve the model’s ability to extract linguistic features from Persian and English inputs, we initially fine-tuned the XLSR for bilingual speech recognition in Persian and English. During this phase, 30% of the training subset of Common Voice Farsi and LibriSpeech (train-clean-100) were used, and the model was trained using CTC loss (Graves et al., 2006) for 40 epochs, with an initial learning rate of 0.001 and an effective batch size of 32. In our experiments, this phase contributes to improving the performance of the phrase verification system.
Finally, to train the classifier, an attention-based pooling layer was added to the fine-tuned XLSR to compute fixed-dimensional utterance-level feature vectors from frame-level representations $h_{t}$ (t=1,,Tt=1,...,T):
@@
where, $e_{t}$ and $\alpha_{t}$ are the attention score and weight, respectively. $\tilde{h}$ refers to the utterance-level feature vector, which is finally fed to a fully connected layer with ReLU activation, followed by a linear classifier. The network was trained using the Cross-Entropy loss function for one epoch on the entire training samples of the challenge dataset, with a learning rate of 0.0005 and an effective batch size of 64.
System Full Training Domain Adaptation Epoch BS LR Epoch BS LR S2 - - - 15 32 3e-4 S3 100 64 1e-3 15 32 3e-4 S4 15 64 1e-3 7 28 5e-5 S5 15 64 1e-3 7 28 5e-5
Table 1: Hyper-parameters used in different submitted systems S2–S5 (BS = batch size, LR = learning rate).
[…]
Table 2: Results of different submissions on the development and evaluation sets.
Subset MinDCF0.01 EER(%) Development 0.0000 0.00 Evaluation 0.0003 0.01
Table 3: Phrase verification performance on TC-vs-TW trials.
@@ 4 Speaker Verification System
To leverage the full power of pre-trained SV models and mitigate the computational cost of training randomly initialized models, we explored two directions for developing our SV system. In the first approach, we fine-tuned several pre-trained ResNet-based models, widely used as a standard architecture in speaker verification. In the second approach, we studied the performance of pre-trained ASR models adapted for SV, which have shown promising results in previous studies (Zhang et al., 2022; Cai et al., 2023; Liao et al., 2023). More specifically, we employed the Whisper-PMFA (Zhao et al., 2024) method, which involves fine-tuning a pre-trained Whisper model for speaker recognition.
@@ 4.1 Training Protocol
  • Full training ($T_{1}$): In this stage, models were trained on a combination of out-of-domain data (Common Voice Farsi and VoxCeleb 1) and in-domain (DeepMine) data, totaling 3,684 speakers, to learn robust and generalizable speaker embeddings across different domains. Pre-trained ResNets did not undergo this stage, as they are already capable of extracting rich speaker-specific features. During this phase, 300 consecutive frames of each input utterance were randomly selected in each epoch to prevent overfitting, reduce GPU memory usage, and accelerate training. Moreover, all augmentation methods explained in Section 2.1 were applied. We employed the widely used AAM-Softmax (Deng et al., 2019) loss with the subcenter method and the Inter-TopK penalty (Zhao et al., 2021) to train our models, with a constant margin and scale of 0.2 and 32, respectively.
  • Domain adaptation ($T_{2}$): We fine-tuned our models using in-domain data after full training to bridge the domain gap and improve performance. During this stage, augmentation methods and the Inter-TopK penalty were removed to prevent domain mismatch. Additionally, the number of randomly selected frames was increased from 300 to 600 to enhance the models’ generalization capability (Garcia-Romero et al., 2019, 2020). Fine-tuning was performed with smaller learning rates to preserve the models’ generalization abilities.
All models were optimized using SGD with a momentum of 0.9 and a weight decay of 1e-4. We also utilized an exponential decay scheduler with a minimum learning rate of 5e-5 for $T_{1}$ and 1e-6 for $T_{2}$. Other training hyper-parameters are listed in Table 1. Note that gradient accumulation was used to achieve the target effective batch size when GPU memory was limited. The dimensionality of speaker embeddings was set to 256 across all models. All experiments were conducted on a single Nvidia RTX 3090 GPU using the WeSpeaker toolkit (Wang et al., 2024).
@@ 4.2 ResNet
ResNet (Xie et al., 2019) is a widely used architecture for speaker recognition that has performed excellently in previous speaker verification challenges (Zheng et al., 2023). Consequently, many open-source implementations and pre-trained models have been publicly released based on this architecture. Trained on large-scale datasets like VoxCeleb 1&2, these pre-trained models can provide a robust starting point for training speaker recognition models on other datasets by improving their generalization and speeding up the convergence.
During the challenge period, we submitted three systems based on a bottleneck-block ResNet, all adopting temporal statistics pooling (Snyder et al., 2018) for aggregating variable-length sequence features into utterance-level embeddings. The first system (S1) was a pre-trained ResNet34 without domain adaptation, while the second one (S2) was a pre-trained ResNet293 that underwent domain adaptation. Finally, we applied both training stages to a randomly initialized ResNet152 to obtain our last ResNet-based system (S3).